# Algebraic Control: Complete Stable Inversion with Necessary and Sufficient Conditions

### v2

@@ III-D Robustness
Note that, the inner product for any vector-valued functions $F,G\in\mathcal{L}_{2}(j\mathbb{R})\cap\mathbb{R}(s)^{n_{y}}$ is defined as:
@@
Let $\operatorname*{Im}_{\mathbb{R}(s)}(P)$ is spanned by an orthonormal basis $\{q_{k}\}_{k=1}^{r}\subset\mathcal{L}_{2}(j\mathbb{R})$. For any $F\in\mathcal{L}_{2}(j\mathbb{R})$, the projection onto $\operatorname*{Im}_{\mathbb{R}(s)}(P)$ is given by:
@@ IV Numerical Examples
In this case, $\operatorname*{rank}_{\mathbb{R}(s)}(P)=2=n_{u}=n_{y}$, which implies that $\operatorname*{res}[Y]=0$. For the desired loop shape $\text{diag}\left[\frac{5}{s},\frac{5}{s}\right]$, we solve (III-C) with $\gamma=3.6$. The approximate inverse, given by $\Xi_{a}^{\Gamma}(s)=S_{i}(s)K(s)$, is shown in Fig. 2.c, consistent with Theorem 3. The resulting control input, $u_{inv}(t)$, is depicted in Fig. 2.b. Compared to the results obtained using the Inner-Outer Factorization (refer to [9, Fig. 8, black lines]), the proposed method significantly improves tracking performance, even using a purely feedforward control scenario.
@@
Figure 2: The output tracking of Example 1 - (a) - corresponding inputs as Theorem 3 - (b) - approximate inversion - (c) -
@@
where $\text{Z}_{P_{\Pi}}=\{-20.4,-6.4,-2.7,-1.2,0.2\}\neq\text{Z}_{P}$ and $\text{Z}_{R_{\Pi}}=\{0.2\}\neq\text{Z}_{R}$. Moreover, $P_{\Pi}(s)$ has poles at $s_{1,2,3,4,5,6}=-10,-10,-4,-4,-0.1,-2.1$, which are different from the poles of the nominal system $P(s)$, $s_{1,2,3,4}=-1,-1,-0.5,-3$. According to Theorem 5, or more specifically, Corollary 5.II, to handle the uncertainty, we employ Fig. 1.(b) with $S_{i}(s)K(s)$ by letting $\epsilon_{i}\to 0$ for both inversion blocks. Two simulation, at $t=125^{th}$ sec, a unit step disturbance applies to the output are run. One is with zero initial states for both inverse and forward systems, and the other one tests the proposed method under random initial states as: $x_{0}=[3.1,4.1,-3.7,4.1,1.3,-4],\bar{x}_{0_{1}}=[-0.2,0.6,1.9,1.9,-0.5,1.9,1.8,0.4,1.4,-0.6,0.3,1.7,1.4,1.8],\bar{x}_{0_{2}}=[0.3,-0.9,0.7,0.9,0.4,0.5,0.5,0.5,-0.2,0.3,-0.7,0.4,-1,-0.4,-0.9].$ The reference tracking performance shown by Fig. 3 with a bounded input validating the theory.
@@
Figure 4: The output tracking of Example 2 - (a) - corresponding inputs as Theorem 4 - (b) - approximate inversion $T_{p}(s)$ (c)
@@ V Conclusion
In this paper, we present a comprehensive approach to stable model inversion aiming to establish necessary and sufficient conditions for stable model inversion using algebraic control.
The proposed approach in this paper applies to various system types, including square, non-square, rank-deficient, and non-minimum phase systems. It is validated by numerical examples is discussed. Our method’s robustness against uncertainties encountered in practical applications expands the model inversion framework and surpasses limitations in existing methods.

### v3

@@ III-D Robustness
Note that the inner product for any vector-valued functions $F,G\in\mathcal{L}_{2}(j\mathbb{R})\cap\mathbb{R}(s)^{n_{y}}$ is defined as:
@@
Let $\operatorname*{Im}_{\mathbb{R}(s)}(P)$ be spanned by an orthonormal basis $\{q_{k}\}_{k=1}^{r}\subset\mathcal{L}_{2}(j\mathbb{R})$. For any $F\in\mathcal{L}_{2}(j\mathbb{R})$, the projection onto $\operatorname*{Im}_{\mathbb{R}(s)}(P)$ is given by:
@@ IV Numerical Examples
In this case, $\operatorname*{rank}_{\mathbb{R}(s)}(P)=2=n_{u}=n_{y}$, which implies that $\operatorname*{res}[Y]=0$. For the desired loop shape $\text{diag}\left[\frac{5}{s},\frac{5}{s}\right]$, we solve (III-C) with $\gamma=3.6$. The approximate inverse $\Xi_{a}^{\Gamma}(s)=S_{i}(s)K(s)$ is plotted in Fig. 2 (right), in line with Theorem 3. The control signal $u_{\text{inv}}(t)$ appears in Fig. 2 (middle), and the output response in Fig. 2 (left). Relative to the inner–outer-factorization benchmark [9, Fig. 8, black curves], our feed-forward design achieves markedly better tracking.
@@
Figure 2: Output tracking (left), control input (middle), and singular-value plot of the approximate inverse $S_{i}(s)K(s)$ (right).
@@
where $\text{Z}_{P_{\Pi}}=\{-20.4,-6.4,-2.7,-1.2,0.2\}\neq\text{Z}_{P}$ and $\text{Z}_{R_{\Pi}}=\{0.2\}\neq\text{Z}_{R}$. Moreover, $P_{\Pi}(s)$ has poles at $s_{1,2,3,4,5,6}=-10,-10,-4,-4,-0.1,-2.1$, which are different from the poles of the nominal system $P(s)$, $s_{1,2,3,4}=-1,-1,-0.5,-3$. According to Theorem 5, or more specifically, Corollary 5.II, to handle the uncertainty, we employ Fig. 1.(b) with $S_{i}(s)K(s)$ by letting $\epsilon_{i}\to 0$ for both inversion blocks. Two simulations were performed. In each, a unit step disturbance is applied to the output at $t=125$ s. One is with zero initial states for both inverse and forward systems, and the other one tests the proposed method under random initial states as: $x_{0}=[3.1,4.1,-3.7,4.1,1.3,-4],\bar{x}_{0_{1}}=[-0.2,0.6,1.9,1.9,-0.5,1.9,1.8,0.4,1.4,-0.6,0.3,1.7,1.4,1.8],\bar{x}_{0_{2}}=[0.3,-0.9,0.7,0.9,0.4,0.5,0.5,0.5,-0.2,0.3,-0.7,0.4,-1,-0.4,-0.9].$ The reference tracking performance shown by Fig. 3 with a bounded input validating the theory.
@@
Figure 4: Output tracking (left), control input (middle), and singular value plot of the approximate inverse $T_{p}(s)$ (right).
@@ V Conclusion
We presented a unified algebraic framework for stable inversion, covering non-minimum-phase, nonsquare, and singular MIMO systems. Necessary and sufficient conditions were established, and constructive inverses were shown with exponential error decay without preview. Uncertainty is handled by an orthogonal projection that isolates the reachable portion of the (desired) output, with the residual captured by the irreducible term $\varphi$. A feed-forward/feedback loop then stabilises the system under uncertainties. Future work will develop data-driven techniques to identify the system’s reachable output subspace and adapt the inversion scheme accordingly.